OpenStudy (anonymous):

how do i find the exact value under the given conditions. sin alpha=20/29, 0

OpenStudy (anonymous):

can anybody help with this????

OpenStudy (jdoe0001):

\(\bf cos(\alpha+\beta)=cos(\alpha)cos(\beta)-sin(\alpha)sin(\beta)\\ \quad \\ sin(\beta)=\cfrac{20}{29}\implies \cfrac{opposite}{hypotenuse}\implies \cfrac{b}{c}\\ \quad \\ c^2=a^2+b^2\implies \sqrt{c^2-b^2}=a\qquad cos(\alpha)=\cfrac{a}{c}\\ \quad \\ \quad \\ -----------------------\\ cos(\beta)=\cfrac{12}{13}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a}{c}\\ \quad \\ c^2=a^2+b^2\implies \sqrt{c^2-a^2}=b\qquad sin(\beta)=\cfrac{b}{c}\)

OpenStudy (jdoe0001):

well... ahemm the sine above . should be alpha but anyhow \(\bf cos(\alpha+\beta)=cos(\alpha)cos(\beta)-sin(\alpha)sin(\beta)\\ \quad \\ sin(\alpha)=\cfrac{20}{29}\implies \cfrac{opposite}{hypotenuse}\implies \cfrac{b}{c}\\ \quad \\ c^2=a^2+b^2\implies \sqrt{c^2-b^2}=a\qquad cos(\alpha)=\cfrac{a}{c}\\ \quad \\ \quad \\ -----------------------\\ cos(\beta)=\cfrac{12}{13}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a}{c}\\ \quad \\ c^2=a^2+b^2\implies \sqrt{c^2-a^2}=b\qquad sin(\beta)=\cfrac{b}{c}\) you'd use the pythagorean theorem to find the missing "leg" or side and get the other trig function, then plug them in the cosine angle sum identity

OpenStudy (anonymous):

got it i will try to figure that out

OpenStudy (anonymous):

thanks i figured the right answer out